# Photonic entanglement enhanced multidimensional spectroscopy for probing exciton correlations: theory and applications to photosynthetic aggregates

### v1

@@ Abstract
Nonlinear spectroscopic techniques using entangled photon pairs provide an opportunity to exploit non-classical correlations encoded in two-photon wavefunctions to manipulate two-exciton wavefunctions. We propose an entangled photon pair-enhanced multidimensional spectroscopic technique which is sensitive to exciton-exciton interactions and correlations at the ultrafast timescale. Simulations for a dissipative photosynthetic aggregate reveal the superior ability of entangled photon pairs, compared to both transform-limited and frequency-chirped laser pulses, to manipulate excited-state absorption pathways. The corresponding spectral features in the two-dimensional spectrogram are interpreted in terms of one- and two-exciton resonances. The signal scales linearly with the incoming intensity of the photon sources. It is argued that classifying these resonances using entangled photon source at the perturbative limit allow for probing exciton correlations at the natural energy scale. These insights can be used to explore multi-exciton dynamics using multiphoton entanglement.
@@ I Introduction
In molecular aggregates, spectroscopic characterization of exciton-exciton interactions and correlations yields critical insights into underlying aggregation properties. However, the range of spectroscopic techniques available for exploring their temporal and spectral signatures, particularly in the presence of phonon-induced dissipation, remains limited. These challenges stem, in part, from a lack of techniques specifically tailored to extract correlation induced features hidden in the dynamics of spatially delocalized two-exciton states. Double-quantum coherence (DQC), a nonlinear spectroscopic scheme, was proposed as a viable approach for probing quasiparticle correlation and their dynamical signatures [40, 33, 41, 39]. Among the suite of nonlinear spectroscopic techniques classified as four-wave mixing, photon echo and transient grating primarily investigate ultrafast exciton transport. Although the observed signals may contain contributions from two-exciton states, the associated resonances are often obscured by dominant one-exciton resonances. The double-quantum coherence technique investigates spectral features specifically associated with two-exciton states; the observed signal is sensitive to the differential nature of the two-exciton wavefunction compared to that of the constituent two one-exciton wavefunctions [9, 54]. In previous investigations, double-quantum coherence spectroscopy has aided the exploration of quasiparticle correlations in molecular aggregates [33, 1], quantum-confined nanostructures [43, 54], correlated lattice systems [6], atomic gases [10, 29, 37], coupled valence and core-hole excitations in molecules [31], vibrations [24], exciton and vibrational polaritons [15, 16, 50, 19].
@@
In this article, we deploy the DQC technique to characterize the ultrafast dynamics of molecular exciton aggregates, specifically identifying the resonances that originate from exciton-exciton interactions and correlations. DQC probes the ultrafast dynamics of the two-exciton wavefunctions via interfering contributions of excited-state absorption (ESA) pathways. Primary signatures of exciton correlations are typically hidden in the two-exciton states, making the technique an ideal candidate to monitor and classify correlation-induced resonances.
@@
For the simulation of the signal, we focus on the light-harvesting complex, LHCII, which acts as the primary photon energy harvester in photosynthetic aggregates. LHCII hosts a large number of one- and two-exciton states delocalized over sites, each of which undergo ultrafast dephasing due to interaction with phonons. As a result, the spectroscopic identification of correlation-induced resonances become particularly challenging. In principle, pulse-shaping techniques are expected to address two-exciton resonances and selectively amplify desired signal components. However, given the relatively high density of optical transitions within a narrow spectral range, classical laser sources face significant challenges. These challenges stem from the unavoidable time-frequency bandwidth constraints. The situation becomes even more critical when identical exciton transitions possess comparable dephasing rates. The entangled photon sources promise to partially alleviate some of these obstacles [3, 32, 27, 28, 26, 21, 22, 49, 15, 16, 14], even though identifying a suitable parameter regime where entangled photon pairs outperform their classical counterparts remains a challenging endeavor [34, 47].
@@
Figure 1: (A) Simplified schematic description of two interacting exciton levels (denoted $|j\rangle$, where $j\in\{a,b,c\})$ at sites $m,n$) highlighting exciton nonlinearities (denoted $\Delta_{m},\Delta_{n}$). The simulation contains $14$ such sites. In the delocalized basis, the two-exciton states are affected by both overtone and combination nonlinearities. (B) The exciton pathways that contribute to the signal are presented using Albrecht notation. The process, involving $\omega_{1}$ and $\omega_{2}$, is common to both pathways. The process, involving $\omega_{3}$ and $\omega_{4}$, is discriminative depending on the differential nature of the two pathways. (C) The corresponding Keldysh-Schwinger diagram is presented.
@@
In what follows, we introduce the exciton-phonon Hamiltonian of the monomeric LHCII, define the one- and two-exciton states, and derive the relevant Green’s functions in section II.1. Subsequently, in section III.2, which is divided into three subsections, we present the analytical signal and corresponding numerical simulations. Simulations utilize entangled photon sources, classical laser pulses, and a variation of the exciton nonlinearity parameters. The article concludes in section IV with a summary and an outlook.
@@ II.1 Hamiltonian
where $B^{\dagger}_{m}(B_{m})$ represent the exciton creation (annihilation) operators, which follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$ (where $\eta=3/2$). Parameters $E_{m}$, $J_{mn}$, $U_{m}^{(1)}$, and $U_{mn}^{(2)}$ represents one-exciton energy, and inter-site hopping, overtone, and combination exciton non-linearity parameters, respectively. $H_{mn}^{(1)}=\sum_{mn}(E_{m}\delta_{mn}+J_{mn})$ denotes the one-exciton Hamiltonian matrix elements, while $H_{mnkl}^{(2)}=H_{mn}^{(1)}\delta_{kl}+\delta_{mn}H_{kl}^{(1)}+U_{m}^{(1)}+U_{mn}^{(2)}$ denote their two-exciton counterparts. While, all parameters $J_{mn}$, $U_{m}^{(1)}$ and $U_{mn}^{(2)}$ affect the overtone and combination two-exciton states, the latter two are responsible for the shift in two-exciton energies. It can be expressed as $E_{m}^{(2)}=2E_{m}^{(2)}+\Delta_{m}$ (where $\Delta_{m}$ is negative is the two-exciton energy is stabilized). In the dipole approximation for the charge densities, $J_{mn}$ are determined by transition dipole moments and $U_{m}^{(1)}$, $U_{mn}^{(2)}$ by the difference of the permanent dipole moments between the excited state and the ground state. Previous theoretical investigations involving two-exciton states which uses coupled two-level models have omitted the role of the overtone states. While arguments based on the laser bandwidth, may allow one to approximate the two-exciton manifold, it would amount to neglecting the role of permanent dipole moments altogether. For chlorophyll molecules, the key pigment in LHCII, the excited state permanent dipole moment is often comparable to their transition dipole moment.
@@
The Hamiltonian yields $N_{g}=1$ ground state, $N_{e}=14$ one-exciton states, and $N_{f}=105$ two-exciton states; the latter is composed of 14 overtone and 91 combination states. The exciton part of the Hamiltonian, given in the first two lines of Eq. (1), can be diagonalized to yield exciton-number-conserving manifolds, denoted as $|g\rangle$, $|e\rangle$, and $|f\rangle$. These states are expressed as $|e_{j}\rangle=\sum_{m}T^{(1)}_{j,m}B_{m}^{\dagger}|0\rangle$ and $|f_{k}\rangle=\sum_{m,n}T^{(2)}_{k,mn}B_{m}^{\dagger}B_{n}^{\dagger}|0\rangle$. The transformation matrix elements, $T_{j,m}^{(1)}$ and $T_{k,mn}^{(2)}$, determine the degree of exciton delocalization for a given exciton state. The exciton states are coupled to collective vibrational modes—the normal modes of intermolecular vibrations—hereafter referred to as phonons. The bare phonon and exciton-phonon coupling Hamiltonians are defined in the third line of Eq. (1). Here, the phonon-mode creation (annihilation) operators $b_{j}^{\dagger}$ ($b_{j}$) follow free-boson commutation relations, defined as $b_{j},b_{i}^{\dagger}]=\delta_{ij}$, while $g_{m,j}$ represents the site-specific coupling strengths. For LHCII, the phonon spectral density function comprises $N_{b_{2}}=48$ high-frequency Brownian oscillator modes and $N_{b_{1}}=1$ low-frequency, overdamped Brownian oscillator mode, expressed as:
@@
The photon-exciton interaction Hamiltonian is given by
$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{V}}$
@@
This dephasing properties are determined by the resonant phonon modes.
@@ III.1 Theory
The DQC signal is related to a specific component of the excitonic nonlinear polarization, $k_{\text{s}}=+k_{1}+k_{2}-k_{3}$ generated as a result of four exciton-photon interactions (see Fig. 1). Following the derivation outlined in the appendix, the final expression is obtained as
@@
where the proportionality constant $s_{0}$ incorporates all the coefficients in the expression and serves as a scaling parameter in the numerical simulation. In the expression, the dipolar spectral weights associated with the two pathways are given by: $w^{(1)}=d_{e^{\prime}f}d_{e^{\prime}g}d_{fe}d_{eg}$ and $w^{(2)}=d_{ge^{\prime}}d_{e^{\prime}f}d_{fe}d_{eg}$. We also introduced two multidimensional functions of the following form
@@
The variables, $\Omega_{j}$, can be scanned to identify and track excitonic resonances. In a typical two-dimensional spectrum (plotted with $\Omega_{2}$ and $\Omega_{3}$ as parametric axes for a fixed $\Omega_{1}$), the off-diagonal spectral features primarily correspond to the spectral signatures that having its origin in exciton interactions and correlations. The four-point photon correlation correlation functions in the numerator allow the modulation of respective spectral weights.
Spontaneous parametric down-conversion (SPDC) provides a control knob for external manipulation of the two-photon wavefunction and the correlation function [46, 11]. In the weak down-conversion regime, the photon correlation function is given by: $\langle E^{{\dagger}}(\omega_{4})E^{{\dagger}}(\omega_{4})E(\omega_{2})E(\omega_{1})\rangle=f^{*}(\omega_{4},\omega_{3})f(\omega_{2},\omega_{1})$, where we define the joint spectral amplitude (JSA) of the two-photon field
@@
$A_{p}(\omega_{a},\omega_{b})$ is the transform-limited field profile of the SPDC pump taken as $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width. The parameter $\alpha$ governs the down-conversion efficiency. The phase function $\phi(\omega_{j},\omega_{k})=(\omega_{j}-\omega_{p}/2)T_{1}+(\omega_{k}-\omega_{p}/2)T_{2}$ is governed by the entanglement time parameter $T_{\text{ent}}=|T_{2}-T_{1}|$. The latter is crucial for spectroscopic applications and is determined by the SPDC crystal length and the relative group velocities of the photons within the crystal. The square modulus of the JSA, depicted in Fig. 6, shows that for strongly anti-correlated pairs, there is parametric freedom to select two distant frequency values. During ultrafast measurements, this flexibility imposes less stringent conditions on the entangled photon pair frequencies. The associated singular values, also presented in Fig. 6, can be used to estimate the effective correlation between spectral modes [34, 22, 7]. The four-point field correlation function for classical laser pulses factorizes into the product of individual amplitude functions, given by: $\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle=A^{*}_{p_{4}}(\omega_{4})A^{*}_{p_{3}}(\omega_{3})A_{p_{2}}(\omega_{2})A_{p_{1}}(\omega_{1})$. The signal, therefore, scales quadratically with the intensity of the incident field. The profile for the linearly chirped pulse is defined as:
$A_{p_{i}}(\omega)$ $=E_{0}\sqrt{\pi/\Gamma_{0,i}}\exp{[-(\omega-\omega_{0,i})^{2}/4\Gamma_{i}]}$ (9)
Here, $1/\Gamma=1/\Gamma_{0}-2i\phi_{0}^{(2)}$, where $\phi_{0}^{(2)}$ denotes the linear chirp rate and $\Gamma_{0}=2\ln(2)/\tau_{0}$ is determined by the temporal pulse width, $\tau_{0}$ [5, 18, 13]. Setting the chirp parameter to zero ($\phi_{0}^{(2)}=0$) yields a transform-limited laser pulse.
@@
Figure 2: Simulated DQC spectra using entangled photon pairs. The sum frequencies of the photon pairs are tuned to the two-exciton states $f_{39}$ (top panel) and $f_{81}$ (bottom panel). The left column provides a reference simulation with an entanglement time $\tilde{T}_{\text{ent}}=10$ fs and an SPDC pump temporal width $\tau_{0}=50$ fs. The middle column shows the effects of varying the SPDC pump temporal width, while the right column shows the effects of varying the entanglement time relative to the references. For a detailed discussion, see Section III.2.1.
@@
The generation of the DQC signal can be conceptualized as a combination of two processes: the initial formation of a state-specific two-exciton coherence, followed by the subsequent projection of these coherent oscillations onto one of two possible one-exciton coherence components. Each of the two pathways depicted in Fig. 1 represents this phenomenology, wherein the two-exciton coherence is generated during the interval $s_{1}+s_{2}$, and the distinct one-exciton coherence are selected during the final interval $s_{3}$. We can manipulate the corresponding two-exciton resonances, $z_{fg}$, by modulating the sum frequency $\omega_{p}=\omega_{1}+\omega_{2}$ during the initial pair of exciton-photon interactions. We can also manipulate the resonances at either $z_{e^{\prime}g}$ or $z_{fe^{\prime}}$ during the subsequent pair of exciton-photon interactions by modulating $\omega_{3}$ and $\omega_{4}$. The final signal is the resultant of two interfering pathways, the magnitude of which is largely determined by these last two interactions. If the contributions from exciton nonlinearities are not significant, the resonances $z_{e^{\prime}g}$ and $z_{fe^{\prime}}$ yield nearly identical contributions. Under these conditions, $z_{fe^{\prime}}\approx\omega_{e^{\prime}g}+\Delta+i\Gamma_{fe^{\prime}}\approx\omega_{e^{\prime}g}+i\Gamma_{fe^{\prime}}$, where the energetic deviation $\Delta$ is negligible. Consequently, the two pathways interfere destructively, and the spectral weights diminish toward zero.
@@
Figure 3: Simulated DQC spectra using classical laser pulses. The central frequencies match the parameters employed for the entangled photon pairs in Fig. 2. The left column presents reference simulations with a temporal width of $\tau_{j,0}=10$ fs. The middle column shows the effect of an increased temporal width ($\tau_{j,0}=20$ fs), which decreases the spectral bandwidth and suppresses several resonances. The right column displays spectra obtained using linearly chirped pulses; notably, the negatively chirped pulse was unable to recover the suppressed resonances. For a detailed discussion, see Section III.2.2.
@@
Figure 4: Simulated DQC spectra using a variation of excitonic nonlinearity parameters which influences both the two-exciton wavefunction and the transition dipole moments. The lower (upper) panel presents the case for classical laser sources (entangled photon pairs). In the upper-left and upper-middle panel, using only higher overtone excitonic nonlinearities and two different two-exciton states are targeted using the same photonic parameter as in the reference simulation in Fig. 2. The upper-right uses degenerate entangled photon parameters to track the down-shifted resonances. The lower-left and lower-middle panel uses higher and lower overtone excitonic nonlinearities(compared to the results in Fig. 3) respectively. The lower-right employs degenerate pulse parameters to explore the down-shifted resonances. For detailed discussions, see Section III.2.3.

### v2

# Photon entanglement-enhanced multidimensional spectroscopy of exciton correlations in photosynthetic aggregates

@@ Abstract
Nonlinear spectroscopic techniques using entangled photon pairs can provide an opportunity to exploit non-classical correlations encoded in two-photon wavefunctions to manipulate two-exciton wavefunctions. We propose an entangled photon pair-enhanced multidimensional spectroscopic technique that is sensitive to exciton-exciton interactions and correlations at the femtosecond timescale. Simulations for a dissipative system, namely, the photosynthetic aggregate reveal the superior ability of entangled photon pairs, compared to both transform-limited and frequency-chirped laser pulses, to manipulate excited-state absorption pathways. The corresponding spectral features in the two-dimensional spectrogram are interpreted in terms of one- and two-exciton resonances. The signal scales linearly with the incoming intensity of the photon sources. We show that classifying these resonances using entangled photon source at the perturbative limit allow for probing exciton correlations at the natural energy scale. These insights can be used to explore multi-exciton dynamics in molecular systems using multiphoton entanglement.
@@ I Introduction
In molecular aggregates, spectroscopic characterization of exciton-exciton interactions and correlations yields critical insights into underlying aggregation properties. However, the range of spectroscopic techniques available for exploring their temporal and spectral signatures, particularly in the presence of phonon-induced dissipation, remains limited. These challenges stem, in part, from a lack of techniques specifically tailored to extract correlation induced features hidden in the dynamics of spatially delocalized two-exciton states. Double-quantum coherence (DQC), a nonlinear spectroscopic scheme, was proposed as a viable approach for probing quasiparticle correlation and its dynamical signatures [40, 33, 41, 39]. Among four-wave mixing techniques, photon echo and transient grating primarily investigate ultrafast exciton transport. Although the observed signals may contain contributions from two-exciton states, the associated resonances are often obscured by dominant one-exciton resonances. The double-quantum coherence technique investigates spectral features specifically associated with two-exciton states; the observed signal is sensitive to the differential nature of the two-exciton wavefunction compared to that of the constituent two one-exciton wavefunctions [9, 54]. In previous investigations, double-quantum coherence spectroscopy has aided the exploration of quasiparticle correlations in molecular aggregates [33, 1], quantum-confined nanostructures [43, 54], correlated lattice systems [6], atomic gases [10, 29, 37], coupled valence and core-hole excitations in molecules [31], vibrations [24], exciton and vibrational polaritons [15, 16, 50, 19].
@@
In this theoretical investigation, we deploy the DQC technique to characterize the ultrafast dynamics of molecular exciton aggregates, specifically identifying the resonances that originate from exciton-exciton interactions and correlations. DQC probes the ultrafast dynamics of the two-exciton wavefunctions via interfering contributions of excited-state absorption (ESA) pathways. Signatures of exciton correlations are typically hidden in the two-exciton states, making the technique the most suitable for monitoring and classifying correlation-induced resonances.
@@
We perform numerical simulations for the light-harvesting complex, LHCII, which acts as the primary photon energy harvester in photosynthetic aggregates. LHCII hosts a large number of one- and two-exciton states delocalized over sites, each of which undergo ultrafast dephasing due to interaction with phonons. As a result, the spectroscopic identification of correlation-induced resonances become particularly challenging. In principle, pulse-shaping techniques are expected to address two-exciton resonances and selectively amplify desired components of the signal. However, given the relatively high density of optical transitions within a narrow spectral range, classical laser sources often face challenges stemming from the unavoidable time-frequency bandwidth constraints. The situation becomes even more critical when identical exciton transitions possess comparable dephasing rates. The entangled photon sources promise to partially alleviate some of these obstacles [3, 32, 27, 28, 26, 21, 22, 49, 15, 16], even though identifying a suitable parameter regime where entangled photon pairs outperform their classical counterparts remains a challenging task [34, 47].
@@
Figure 1: (A) Schematic representation of two interacting excitonic sites, $m$ and $n$. Each site comprises three energy levels, $|j\rangle$, where $j\in\{a,b,c\})$, highlighting overtone exciton nonlinearities $\Delta_{m}$ and $\Delta_{n}$. The model includes 14 such sites. In the delocalized basis, the two-exciton states $|f\rangle$ are affected by both overtone and combination nonlinearities. (B) Exciton pathways contributing to the signal, shown in Albrecht notation. Processes involving $\omega_{1}$ and $\omega_{2}$ are common to both pathways. Processes involving $\omega_{3}$ and $\omega_{4}$ determine the differences between the two pathways and are, therefore, discriminative. (C) Corresponding Keldysh-Schwinger loop diagrams for the two pathways contributing to the signal.
@@
In what follows, we introduce the exciton-phonon Hamiltonian of the monomeric LHCII, define the one- and two-exciton states, and derive the relevant Green’s functions in section II.1. In section III.2, which is divided into three subsections, we present the analytical signal and corresponding numerical simulations. Simulations utilize entangled photon sources, classical laser pulses, and a variation of the exciton nonlinearity parameters. The article concludes in section IV with a summary and an outlook.
@@ II.1 Hamiltonian
where $B^{\dagger}_{m}(B_{m})$ represent the exciton creation (annihilation) operators, which follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$ (where $\eta=3/2$). Parameters $E_{m}$, $J_{mn}$, $U_{m}^{(1)}$, and $U_{mn}^{(2)}$ represents one-exciton energy, and inter-site hopping, overtone, and combination exciton non-linearity parameters, respectively. $H_{mn}^{(1)}=\sum_{mn}(E_{m}\delta_{mn}+J_{mn})$ denotes the one-exciton Hamiltonian matrix elements, while $H_{mnkl}^{(2)}=H_{mn}^{(1)}\delta_{kl}+\delta_{mn}H_{kl}^{(1)}+U_{m}^{(1)}+U_{mn}^{(2)}$ denote their two-exciton counterparts. While, all parameters $J_{mn}$, $U_{m}^{(1)}$ and $U_{mn}^{(2)}$ affect the overtone and combination two-exciton states, the latter two are responsible for the shift in two-exciton energies. It can be expressed as $E_{m}^{(2)}=2E_{m}^{(2)}+\Delta_{m}$ (where $\Delta_{m}$ is negative is the two-exciton energy is stabilized). In the dipole approximation for the charge densities, $J_{mn}$ are determined by transition dipole moments, while $U_{m}^{(1)}$, $U_{mn}^{(2)}$ are determined by the difference between the permanent dipole moments of the excited and ground states. Previous theoretical investigations of two-exciton states based on coupled two-level models have omitted the role of overtone states. Whereas arguments based on laser bandwidth might justify limiting the number of accessible states in the two-exciton manifold, it would amount to neglecting the role of permanent dipole moments. For chlorophyll molecules, the primary pigments in LHCII, the excited-state permanent dipole moment is often similar in magnitude to the transition dipole moment.
@@
The Hamiltonian yields $N_{g}=1$ ground state, $N_{e}=14$ one-exciton states, and $N_{f}=105$ two-exciton states; the latter comprises 14 overtone and 91 combination states. The exciton part of the Hamiltonian, given in the first two lines of Eq. (1), is diagonalized to yield exciton-number-conserving manifolds, denoted as $|g\rangle$, $|e\rangle$, and $|f\rangle$. These states are expressed as $|e_{j}\rangle=\sum_{m}T^{(1)}_{j,m}B_{m}^{\dagger}|0\rangle$ and $|f_{k}\rangle=\sum_{m,n}T^{(2)}_{k,mn}B_{m}^{\dagger}B_{n}^{\dagger}|0\rangle$. The transformation matrix elements, $T_{j,m}^{(1)}$ and $T_{k,mn}^{(2)}$, determine the degree of exciton delocalization for a given exciton state. The exciton states are coupled to collective vibrational modes—the normal modes of intermolecular vibrations—hereafter referred to as phonons. The bare phonon and exciton-phonon coupling Hamiltonians are defined in the third line of Eq. (1). Here, the phonon-mode creation (annihilation) operators $b_{j}^{\dagger}$ ($b_{j}$) follow free-boson commutation relations, defined as $b_{j},b_{i}^{\dagger}]=\delta_{ij}$, while $g_{m,j}$ represents the site-specific coupling strengths. For LHCII, the phonon spectral density function comprises $N_{b_{2}}=48$ high-frequency Brownian oscillator modes and $N_{b_{1}}=1$ low-frequency, overdamped Brownian oscillator mode, expressed as:
@@
The exciton-photon interaction Hamiltonian is given by
$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{V_{j}}}$
@@
The dephasing properties are determined by the resonant phonon modes.
@@ III.1 Theory
The DQC signal is related to a specific component of the excitonic nonlinear polarization, $k_{\text{s}}=+k_{1}+k_{2}-k_{3}$ generated as a result of four exciton-photon interactions (see Fig. 1). Following the derivation outlined in the Appendix B, the final expression is obtained as
@@
The proportionality constant $s_{0}$ incorporates all coefficients in the expression and serves as a scaling parameter for the numerical simulation. The dipolar spectral weights associated with the two pathways are given by: $w^{(1)}=d_{e^{\prime}f}d_{e^{\prime}g}d_{fe}d_{eg}$ and $w^{(2)}=d_{ge^{\prime}}d_{e^{\prime}f}d_{fe}d_{eg}$. We introduced two multidimensional functions
@@
The variables $\Omega_{j}$ are scanned to identify and track excitonic resonances. In a typical two-dimensional spectrum, plotted with $\Omega_{2}$ and $\Omega_{3}$ as the axes for a fixed $\Omega_{1}$, the off-diagonal spectral features primarily represent signatures originating from exciton interactions and correlations. The four-point photon correlation correlation functions in the numerator allow the modulation of respective spectral weights.
Spontaneous parametric down-conversion (SPDC) provides a mechanism for the external manipulation of the two-photon wavefunction and the correlation function [46, 11]. In the weak down-conversion regime, the photon correlation function is given by: $\langle E^{{\dagger}}(\omega_{4})E^{{\dagger}}(\omega_{4})E(\omega_{2})E(\omega_{1})\rangle=f^{*}(\omega_{4},\omega_{3})f(\omega_{2},\omega_{1})$, where we define the joint spectral amplitude (JSA) of the two-photon field
@@
Here, the transform-limited field profile of the SPDC pump is given by $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width. The parameter $\alpha$ governs the down-conversion efficiency. The phase function $\phi(\omega_{j},\omega_{k})=(\omega_{j}-\omega_{p}/2)T_{1}+(\omega_{k}-\omega_{p}/2)T_{2}$ is governed by the entanglement time parameter $T_{\text{ent}}=|T_{2}-T_{1}|$. The latter is crucial for spectroscopic applications. It is determined by the length of the SPDC crystal and the relative group velocities of the photon pairs within the crystal. The square modulus of the JSA, depicted in Fig. 6, shows that for strongly anti-correlated pairs, there is a parametric freedom to select two distant frequency values. In ultrafast measurements, this flexibility imposes less stringent conditions on the frequencies of the entangled photon pairs. The associated singular values, also presented in Fig. 6, are used to estimate the effective correlation between spectral modes [34, 22, 7]. In comparison, the four-point field correlation function for classical laser pulses factorizes into the product of individual amplitude functions: $\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle=A^{*}_{p_{4}}(\omega_{4})A^{*}_{p_{3}}(\omega_{3})A_{p_{2}}(\omega_{2})A_{p_{1}}(\omega_{1})$. The signal, therefore, scales quadratically with the intensity of the incident field. The frequency domain representation of a transform-limited pulse with a Gaussian profile is given by
$A_{p_{j}}(\omega)$ $=E_{0}\sqrt{\pi/\Gamma_{0}}\exp{[-(\omega-\omega_{0,j})^{2}/4\Gamma_{0,j}]}$ (9)
where $\Gamma_{0,j}=1/T_{0,j}^{2}=2\ln 2/\tau_{0,j}^{2}$, $T_{0,j}$ is the temporal half-width at $1/e$, and $\tau_{0,j}$ is the temporal width of the pulse. Pulse shaping techniques effectively adds a frequency dependent phase which, restricted to the second-order dispersion (additionally, neglecting the constant phase and the linear order terms), can be expressed as
$A_{c,p_{j}}(\omega)$ $=E_{0}\sqrt{\pi/\Gamma_{0,j}}\exp{[-(\omega-\omega_{0,j})^{2}/4\Gamma_{j}]}$ (10)
The profile above is used for the linearly chirped pulse. Here, $1/\Gamma_{j}=1/\Gamma_{0,j}-2i\phi_{0,j}^{(2)}$, where $\phi_{0,j}^{(2)}$ is related to the linear chirp rate [5, 18, 13]. We introduce $\Gamma_{j}=1/T_{p,j}^{2}+i\alpha_{j}$, with $T_{p,j}=T_{0,j}\sqrt{1+(2\phi_{0,j}^{(2)}/T_{0,j}^{2})^{2}}$ and $\alpha_{j}=2\phi_{0,j}^{(2)}/[T_{0,j}^{4}+(2\phi_{0,j}^{(2)})^{2}]$. Here, $\alpha$ is the chirp rate satisfying: $\omega_{j}(t)=\omega_{0,j}+2\alpha_{j}t$. Setting the quadratic dispersion to zero, i.e., $\phi_{0,j}^{(2)}=0$ yields transform-limited laser pulse back. In our case, two classes of optical processes induced by the laser pulses, namely absorption and stimulated emissions, involve a large number of excitonic states. The time-dependent instantaneous frequency is suitable for simultaneously controlling a finite number of transitions within the spectral bandwidth.
Conceptually, the generation of the DQC signal can be viewed as a combination of two processes: the initial formation of a state-specific two-exciton coherence, followed by the projection of these coherent oscillations onto one of two possible one-exciton coherence components.
@@
Figure 2: Simulated DQC spectra using entangled photon pairs. The sum frequencies of the photon pairs are tuned to the two-exciton states $f_{39}$ (top panel) and $f_{81}$ (bottom panel). The top-left and bottom-left panels serve as reference simulations with $\tilde{T}_{\text{ent}}=10\,\text{fs}$ and $\tau_{0}=50\,\text{fs}$. The middle and right columns show the effects of varying the temporal width of the SPDC pump and the entanglement time, respectively. The top-middle and bottom-middle panels use $\tau_{0}=100$ fs and $\tau_{0}=20$ fs, respectively, while the top-right and bottom-right panels use $\tilde{T}_{\text{ent}}=60$ fs. All other parameters remain identical to the reference. For a detailed discussion, see Section III.2.1.
@@
The two pathways depicted in Fig. 1 captures this phenomenology. Here, the two-exciton coherence is generated during the interval $s_{1}+s_{2}$, and the one-exciton coherence components are generated during the final interval $s_{3}$. We can manipulate the two-exciton resonances, $z_{fg}$, by modulating the sum frequency $\omega_{p}=\omega_{1}+\omega_{2}$ during the initial pair of exciton-photon interactions. We can also manipulate the one-exciton resonances at $z_{e^{\prime}g}$ or $z_{fe^{\prime}}$, by modulating $\omega_{3}$ and $\omega_{4}$, during the subsequent pair of exciton-photon interactions. The final signal is the resultant of two interfering pathways, the magnitude of which is largely determined by the last two interactions. If the contributions from exciton nonlinearities are not significant, the resonances $z_{e^{\prime}g}$ and $z_{fe^{\prime}}$ yield nearly identical contributions. In such situations, we have $z_{fe^{\prime}}\approx\omega_{e^{\prime}g}+\Delta+i\Gamma_{fe^{\prime}}\approx\omega_{e^{\prime}g}+i\Gamma_{fe^{\prime}}$, where the energetic deviation $\Delta$ is negligible. Consequently, the two pathways interfere destructively, and the spectral weights diminish toward zero.
@@
Figure 3: Simulated DQC spectra using classical laser pulses. The central frequencies of the laser pulses match the frequencies of the entangled photon pairs deployed in Fig. 2. The top- and bottom-left column present the reference simulations with $\tau_{j,0}=10$ fs. The middle column shows the effect of an increased temporal width by using $\tau_{j,0}=20$ fs. The decrease in the spectral bandwidth suppresses several peaks. The right column displays spectra obtained using linearly chirped pulses. The negatively chirped pulse was unable to recover the suppressed peaks. For a detailed discussion, see Section III.2.2.
@@
Figure 4: Simulated DQC spectra evaluated using two distinct sets of excitonic nonlinearity parameters. These parameters influence both the two-exciton wavefunction and the transition dipole moments, as highlighted in the Appendix A. The top and bottom panels present simulations using the entangled photon pairs and classical laser pulses, respectively. The upper panels utilize larger overtone excitonic nonlinearities (values given in the lower panel of the Fig. 5) than previous simulations in Fig. 2. The top-left and top-middle panels employ the same photonic parameters as those in the top-left and bottom-left of the same. The top-right panel uses degenerate entangled photon pairs with frequencies $\omega_{j}=E_{09}$ The bottom-left and bottom-middle panels use larger and smaller overtone excitonic nonlinearities, respectively. The bottom-right panel employs degenerate laser pulses to explore the down-shifted resonances. For a detailed discussion, see Section III.2.3.